Find the equation of the line containing the point (-1,3) and perpendicular to hte line 2x+4y=-1
Answer by fractalier(6550)

We need a line perpendicular to
2x + 4y = -1
Let us find the slope of this line first...
4y = -2x - 1
y = (-1/2)x - 1/4
The slope here is -1/2.
The slope of the perpendicular to it is the negative reciprocal, or +2.
Now use the point-slope form
y - y1 = m(x - x1)
y - 3 = 2(x + 1)
y - 3 = 2x + 2
y = 2x + 5